# Dusit Niyato, , Bo Ai

### v1

@@ Appendix F Proof of Corollary 2
We first reconstruct the SINR in (26). Note that $\sum_{m=1}^{M}{\mathbb{E}\{\mathbf{v}_{mk}^{H}\mathbf{g}_{mk}\}}=\sum_{m=1}^{M}{\mathbf{w}_{mk}^{H}\mathrm{vec}(\mathbb{E}\{\mathbf{g}_{mk}\hat{\mathbf{g}}_{mk}^{H}\})}$ by applying (32) and (34), where $\mathbf{w}_{mk}=\mathrm{vec}(\mathbf{W}_{mk})\in\mathbb{C}^{N^{2}}$. Then, by defining $\mathbf{w}_{k}^{\mathrm{d}}=\left[\mathbf{w}_{1k}^{T},\mathbf{w}_{2k}^{T},\dots,\mathbf{w}_{Mk}^{T}\right]^{T}\in\mathbb{C}^{MN^{2}}$ and $\mathbf{h}_{k}=[\mathrm{vec}(\mathbf{g}_{1k}\hat{\mathbf{g}}_{1k}^{H})^{T},\dots,\mathrm{vec}(\mathbf{g}_{Mk}\hat{\mathbf{g}}_{Mk}^{H})^{T}]^{T}\in\mathbb{C}^{MN^{2}}$, we can further represent it as $\sum_{m=1}^{M}{\mathbb{E}\{\mathbf{v}_{mk}^{H}\mathbf{g}_{mk}\}}=\mathbf{w}_{k}^{\mathrm{d},H}\mathbb{E}\{\mathbf{h}_{k}\}$. As for $\mathbb{E}\{|\sum_{m=1}^{M}{\mathbf{v}_{mk}^{H}\mathbf{g}_{ml}}|^{2}\}$, we have
$\mathbb{E}\{|\sum\nolimits_{m=1}^{M}{\mathbf{v}_{mk}^{H}\mathbf{g}_{ml}}|^{2}\}\overset{(a)}{=}\mathbb{E}\{|\sum\nolimits_{m=1}^{M}{\mathrm{tr}(\mathbf{W}_{mk}^{H}\mathbf{g}_{ml}\hat{\mathbf{g}}_{mk}^{H})}|^{2}\}$ (46)
=$\overset{(b)}{=}\mathbb{E}\{|\sum\nolimits_{m=1}^{M}{\mathbf{w}_{mk}^{H}\mathrm{vec}(\mathbf{g}_{ml}\hat{\mathbf{g}}_{mk}^{H})}|^{2}\}=\mathbb{E}\{|\mathbf{w}_{k}^{\mathrm{d},H}\mathbf{z}_{kl}|^{2}\}$
@@
where step (a) and step (b) follows from (32) and (34), respectively, with $\mathbf{z}_{kl}=[\mathrm{vec}(\mathbf{g}_{1l}\hat{\mathbf{g}}_{1k}^{H})^{T},\dots,\mathrm{vec}(\mathbf{g}_{Ml}\hat{\mathbf{g}}_{Mk}^{H})^{T}]\in\mathbb{C}^{MN^{2}}$. Moreover, we have
$\sum\nolimits_{m=1}^{M}{\mathbb{E}\{\|\mathbf{v}_{mk}\|^{2}\}}{=}\sum\nolimits_{m=1}^{M}{\mathrm{tr}(\mathbf{W}_{mk}^{H}\mathbf{W}_{mk}\mathbb{E}\{\hat{\mathbf{g}}_{mk}\hat{\mathbf{g}}_{mk}^{H}\})}$ (47)
=$\overset{(a)}{=}\sum\nolimits_{m=1}^{M}{\mathbf{w}_{mk}^{H}(\mathbb{E}\{\hat{\mathbf{g}}_{mk}\hat{\mathbf{g}}_{mk}^{H}\}^{T}\otimes\mathbf{I}_{N})}\mathbf{w}_{mk}=\mathbf{w}_{k}^{H}\mathbf{\Theta}_{k}\mathbf{w}_{k},$
where step (a) follows from (36), with $\mathbf{\Theta}_{k}=\mathrm{diag}\{(\mathbb{E}\{\hat{\mathbf{g}}_{1k}\hat{\mathbf{g}}_{1k}^{H}\}^{T}\otimes\mathbf{I}_{N}),\dots,(\mathbb{E}\{\hat{\mathbf{g}}_{Mk}\hat{\mathbf{g}}_{Mk}^{H}\}^{T}\otimes\mathbf{I}_{N})\}\in\mathbb{C}^{MN^{2}\times MN^{2}}$. Combining all above terms, we can rewrite the SINR in (26) as
$\mathrm{SINR}_{k}^{\mathrm{d}}\!\!=\!\!\frac{p_{k}|\mathbf{w}_{k}^{\mathrm{d},H}\mathbb{E}\{\mathbf{h}_{k}\}|^{2}}{\mathbf{w}_{k}^{\mathrm{d},H}\!(\!\sum\limits_{l=1}^{K}{p_{l}\mathbb{E}\{\mathbf{z}_{kl}\mathbf{z}_{kl}^{H}\}\!-\!p_{k}\mathbb{E}\{\mathbf{h}_{k}\}\mathbb{E}\{\mathbf{h}_{k}\}^{H}\!\!+\!\!\sigma^{2}\mathbf{\Theta}_{k}})\mathbf{w}^{\mathrm{d}}_{k}}.$ (48)
Note that $\mathrm{SINR}_{k}^{\mathrm{d}}$ in (48) is a generalized Rayleigh quotient with respect to $\mathbf{w}^{\mathrm{d}}_{k}$. Thus, we can derive the optimal $\mathbf{w}_{k}^{\mathrm{d},*}$, which can maximize $\mathrm{SINR}_{k}^{\mathrm{d}}$ as in (27) with the maximum value $\mathrm{SINR}_{k}^{\mathrm{d},*}=p_{k}\mathbb{E}\{\mathbf{h}_{k}\}^{H}\mathbf{w}_{k}^{\mathrm{d},*}$. Then, by applying the vectorization reversion to $\mathbf{w}_{k}^{\mathrm{d},*}=[\mathbf{w}_{1k}^{*,T},\dots,\mathbf{w}_{Mk}^{*,T}]^{T}\in\mathbb{C}^{MN^{2}}$ for each AP, we can derive the optimal $\mathbf{W}_{mk}^{*}=\mathrm{vec}^{-1}(\mathbf{w}_{mk}^{*})$.
@@ Appendix G Proof of Theorem 4
Based on Theorem 2, we can easily compute (26) in closed-form as (40) on the top of the previous page, where $\gamma_{mkl}=\mathrm{tr}(\mathbf{W}_{mk}^{H}\mathbf{B}_{mlk})$. Then, we have $|\sum_{m=1}^{M}{\mathrm{tr}\left(\mathbf{W}_{mk}^{H}\bar{\mathbf{R}}_{mk}\right)}|^{2}=|\sum_{m=1}^{M}{\mathbf{w}_{mk}^{H}\overline{\mathbf{h}}_{mk}}|^{2}=|\mathbf{w}_{k}^{\mathrm{d},H}\overline{\mathbf{h}}_{k}|^{2}$. Then, by applying results in Appendix A, we can rewrite $\varphi_{mkl}$, $\upsilon_{mkl}$, and $\gamma_{mkl}$ in (40) as $\varphi_{mkl}=\mathbf{w}_{mk}^{H}(\bar{\mathbf{G}}_{mkk}^{T}\otimes\bar{\mathbf{G}}_{mll})\mathbf{w}_{mk}+\mathbf{w}_{mk}^{H}(\bar{\mathbf{G}}_{mkk}^{T}\otimes\check{\mathbf{R}}_{ml})\mathbf{w}_{mk}+\mathbf{w}_{mk}^{H}(\hat{\mathbf{R}}_{mk}^{T}\otimes\bar{\mathbf{G}}_{mll})\mathbf{w}_{mk}+\mathbf{w}_{mk}^{H}(\hat{\mathbf{R}}_{mk}^{T}\otimes\check{\mathbf{R}}_{ml})\mathbf{w}_{mk}$, $\upsilon_{mkl}=p_{k}p_{l}\tau_{p}^{2}|\mathbf{w}_{mk}^{H}\tilde{\mathbf{r}}_{mlk}|^{2}$ for $l\in\mathcal{P}_{k}\backslash\{k\}$ and $p_{k}p_{l}\tau_{p}^{2}|\mathbf{w}_{mk}^{H}\tilde{\mathbf{r}}_{mlk}|+p_{k}^{2}\tau_{p}\bar{\mathbf{g}}_{mkk}\tilde{\mathbf{r}}_{mkk}^{H}+p_{k}^{2}\tau_{p}\tilde{\mathbf{r}}_{mkk}\bar{\mathbf{g}}_{mkk}^{H}$ for $l=k$, and $\gamma_{mkl}=\mathbf{w}_{mk}^{H}\mathbf{b}_{mlk}$, respectively. By defining $\mathbf{w}_{k}^{\mathrm{d}}=\left[\mathbf{w}_{1k}^{T},\mathbf{w}_{2k}^{T},\dots,\mathbf{w}_{Mk}^{T}\right]^{T}\in\mathbb{C}^{MN^{2}}$ and based on above results, we can easily construct the diagonal block matrices $\mathbf{\Lambda}_{k}^{(1)}$, $\mathbf{\Lambda}_{k}^{(2)}$, and $\mathbf{\Lambda}_{k}^{(4)}$. Moreover, we have $|\sum_{m=1}^{M}{\gamma_{mkl}}|^{2}=|\sum_{m=1}^{M}{\mathbf{w}_{mk}^{H}\mathbf{b}_{mlk}}|^{2}=|\mathbf{w}_{k}^{\mathrm{d},H}\mathbf{b}_{lk}|^{2}$ and $\sum\nolimits_{m=1}^{M}{|\gamma_{mkl}|^{2}}=\sum\nolimits_{m=1}^{M}{|\mathbf{w}_{mk}^{H}\mathbf{b}_{mlk}|^{2}=\mathbf{w}_{k}^{\mathrm{d},H}\bar{\mathbf{B}}_{lk}\mathbf{w}_{k}^{\mathrm{d}}}$. Thus, we can derive $\mathbf{\Lambda}_{k}^{(3)}$. Plugging all derived results into (40) and applying the property of the generalized Rayleigh quotient, we can obtain the closed-form results in Theorem 4.

### v2

@@ Appendix F Proof of Corollary 2
We first reconstruct the SINR in (26). Note that $\sum_{m=1}^{M}{\mathbb{E}\{\mathbf{v}_{mk}^{H}\mathbf{g}_{mk}\}}=\sum_{m=1}^{M}{\mathbf{w}_{mk}^{H}\mathrm{vec}(\mathbb{E}\{\mathbf{g}_{mk}\widehat{\mathbf{g}}_{mk}^{H}\})}$ by applying (32) and (34), where $\mathbf{w}_{mk}=\mathrm{vec}(\mathbf{W}_{mk})\in\mathbb{C}^{N^{2}}$. Then, by defining $\mathbf{w}_{k}^{\mathrm{d}}=\left[\mathbf{w}_{1k}^{T},\mathbf{w}_{2k}^{T},\dots,\mathbf{w}_{Mk}^{T}\right]^{T}\in\mathbb{C}^{MN^{2}}$ and $\mathbf{h}_{k}=[\mathrm{vec}(\mathbf{g}_{1k}\widehat{\mathbf{g}}_{1k}^{H})^{T},\dots,\mathrm{vec}(\mathbf{g}_{Mk}\widehat{\mathbf{g}}_{Mk}^{H})^{T}]^{T}\in\mathbb{C}^{MN^{2}}$, we can further represent it as $\sum_{m=1}^{M}{\mathbb{E}\{\mathbf{v}_{mk}^{H}\mathbf{g}_{mk}\}}=\mathbf{w}_{k}^{\mathrm{d},H}\mathbb{E}\{\mathbf{h}_{k}\}$. As for $\mathbb{E}\{|\sum_{m=1}^{M}{\mathbf{v}_{mk}^{H}\mathbf{g}_{ml}}|^{2}\}$, we have
$\mathbb{E}\{|\sum\nolimits_{m=1}^{M}{\mathbf{v}_{mk}^{H}\mathbf{g}_{ml}}|^{2}\}\overset{(a)}{=}\mathbb{E}\{|\sum\nolimits_{m=1}^{M}{\mathrm{tr}(\mathbf{W}_{mk}^{H}\mathbf{g}_{ml}\widehat{\mathbf{g}}_{mk}^{H})}|^{2}\}$ (46)
=$\overset{(b)}{=}\mathbb{E}\{|\sum\nolimits_{m=1}^{M}{\mathbf{w}_{mk}^{H}\mathrm{vec}(\mathbf{g}_{ml}\widehat{\mathbf{g}}_{mk}^{H})}|^{2}\}=\mathbb{E}\{|\mathbf{w}_{k}^{\mathrm{d},H}\mathbf{z}_{kl}|^{2}\}$
@@
where step (a) and step (b) follows from (32) and (34), respectively, with $\mathbf{z}_{kl}=[\mathrm{vec}(\mathbf{g}_{1l}\widehat{\mathbf{g}}_{1k}^{H})^{T},\dots,\mathrm{vec}(\mathbf{g}_{Ml}\widehat{\mathbf{g}}_{Mk}^{H})^{T}]\in\mathbb{C}^{MN^{2}}$. Moreover, we have
$\sum\nolimits_{m=1}^{M}{\mathbb{E}\{\|\mathbf{v}_{mk}\|^{2}\}}{=}\sum\nolimits_{m=1}^{M}{\mathrm{tr}(\mathbf{W}_{mk}^{H}\mathbf{W}_{mk}\mathbb{E}\{\widehat{\mathbf{g}}_{mk}\widehat{\mathbf{g}}_{mk}^{H}\})}$ (47)
=$\overset{(a)}{=}\sum\nolimits_{m=1}^{M}{\mathbf{w}_{mk}^{H}(\mathbb{E}\{\widehat{\mathbf{g}}_{mk}\widehat{\mathbf{g}}_{mk}^{H}\}^{T}\otimes\mathbf{I}_{N})}\mathbf{w}_{mk}=\mathbf{w}_{k}^{H}\mathbf{\Theta}_{k}\mathbf{w}_{k},$
where step (a) follows from (36), with $\mathbf{\Theta}_{k}=\mathrm{diag}\{(\mathbb{E}\{\widehat{\mathbf{g}}_{1k}\widehat{\mathbf{g}}_{1k}^{H}\}^{T}\otimes\mathbf{I}_{N}),\dots,(\mathbb{E}\{\widehat{\mathbf{g}}_{Mk}\widehat{\mathbf{g}}_{Mk}^{H}\}^{T}\otimes\mathbf{I}_{N})\}\in\mathbb{C}^{MN^{2}\times MN^{2}}$. Combining all above terms, we can rewrite the SINR in (26) as
$\widetilde{\mathrm{SINR}}_{k}^{\mathrm{LSFD}}=$ (48)
$\frac{p_{k}|\mathbf{w}_{k}^{\mathrm{d},H}\mathbb{E}\{\mathbf{h}_{k}\}|^{2}}{\mathbf{w}_{k}^{\mathrm{d},H}\!(\!\sum\limits_{l=1}^{K}{p_{l}\mathbb{E}\{\mathbf{z}_{kl}\mathbf{z}_{kl}^{H}\}\!-\!p_{k}\mathbb{E}\{\mathbf{h}_{k}\}\mathbb{E}\{\mathbf{h}_{k}\}^{H}\!\!+\!\!\sigma^{2}\mathbf{\Theta}_{k}})\mathbf{w}^{\mathrm{d}}_{k}}.$
Note that $\mathrm{SINR}_{k}^{\mathrm{d}}$ in (LABEL:Dis_SINR_expansion) is a generalized Rayleigh quotient with respect to $\mathbf{w}^{\mathrm{d}}_{k}$. Thus, we can derive the optimal $\mathbf{w}_{k}^{\mathrm{d},*}$ in (LABEL:DG_OBE_W_Monte), which can maximize $\mathrm{SINR}_{k}^{\mathrm{d}}$ with the maximum value $\mathrm{SINR}_{k}^{\mathrm{d},*}=p_{k}\mathbb{E}\{\mathbf{h}_{k}\}^{H}\mathbf{w}_{k}^{\mathrm{d},*}$. Then, by applying the vectorization reversion to $\mathbf{w}_{k}^{\mathrm{d},*}=[\mathbf{w}_{1k}^{*,T},\dots,\mathbf{w}_{Mk}^{*,T}]^{T}\in\mathbb{C}^{MN^{2}}$ for each AP, we can derive the optimal $\mathbf{W}_{mk}^{*}=\mathrm{vec}^{-1}(\mathbf{w}_{mk}^{*})$.
@@ Appendix G Proof of Theorem 4
Based on Theorem 2, we can easily compute (26) in closed-form as (40) on the top of this page, where $\gamma_{mkl}=\mathrm{tr}(\mathbf{W}_{mk}^{H}\mathbf{B}_{mlk})$. Then, we have $|\sum_{m=1}^{M}{\mathrm{tr}\left(\mathbf{W}_{mk}^{H}\overline{\mathbf{R}}_{mk}\right)}|^{2}=|\sum_{m=1}^{M}{\mathbf{w}_{mk}^{H}\overline{\mathbf{h}}_{mk}}|^{2}=|\mathbf{w}_{k}^{\mathrm{d},H}\overline{\mathbf{h}}_{k}|^{2}$. Then, by applying results in Appendix A, we can rewrite $\varphi_{mkl}$, $\upsilon_{mkl}$, and $\gamma_{mkl}$ in (40) as $\varphi_{mkl}=\mathbf{w}_{mk}^{H}(\overline{\mathbf{G}}_{mkk}^{T}\otimes\overline{\mathbf{G}}_{mll})\mathbf{w}_{mk}+\mathbf{w}_{mk}^{H}(\overline{\mathbf{G}}_{mkk}^{T}\otimes\check{\mathbf{R}}_{ml})\mathbf{w}_{mk}+\mathbf{w}_{mk}^{H}(\widehat{\mathbf{R}}_{mk}^{T}\otimes\overline{\mathbf{G}}_{mll})\mathbf{w}_{mk}+\mathbf{w}_{mk}^{H}(\widehat{\mathbf{R}}_{mk}^{T}\otimes\check{\mathbf{R}}_{ml})\mathbf{w}_{mk}$, $\upsilon_{mkl}=p_{k}p_{l}\tau_{p}^{2}|\mathbf{w}_{mk}^{H}\tilde{\mathbf{r}}_{mlk}|^{2}$ for $l\in\mathcal{P}_{k}\backslash\{k\}$ and $p_{k}p_{l}\tau_{p}^{2}|\mathbf{w}_{mk}^{H}\tilde{\mathbf{r}}_{mlk}|+p_{k}^{2}\tau_{p}\overline{\mathbf{g}}_{mkk}\tilde{\mathbf{r}}_{mkk}^{H}+p_{k}^{2}\tau_{p}\tilde{\mathbf{r}}_{mkk}\overline{\mathbf{g}}_{mkk}^{H}$ for $l=k$, and $\gamma_{mkl}=\mathbf{w}_{mk}^{H}\mathbf{b}_{mlk}$, respectively. By defining $\mathbf{w}_{k}^{\mathrm{d}}=\left[\mathbf{w}_{1k}^{T},\mathbf{w}_{2k}^{T},\dots,\mathbf{w}_{Mk}^{T}\right]^{T}\in\mathbb{C}^{MN^{2}}$ and based on above results, we can easily construct the diagonal block matrices $\mathbf{\Lambda}_{k}^{(1)}$, $\mathbf{\Lambda}_{k}^{(2)}$, and $\mathbf{\Lambda}_{k}^{(4)}$. Moreover, we have $|\sum_{m=1}^{M}{\gamma_{mkl}}|^{2}=|\sum_{m=1}^{M}{\mathbf{w}_{mk}^{H}\mathbf{b}_{mlk}}|^{2}=|\mathbf{w}_{k}^{\mathrm{d},H}\mathbf{b}_{lk}|^{2}$ and $\sum\nolimits_{m=1}^{M}{|\gamma_{mkl}|^{2}}=\sum\nolimits_{m=1}^{M}{|\mathbf{w}_{mk}^{H}\mathbf{b}_{mlk}|^{2}=\mathbf{w}_{k}^{\mathrm{d},H}\overline{\mathbf{B}}_{lk}\mathbf{w}_{k}^{\mathrm{d}}}$. Thus, we can derive $\mathbf{\Lambda}_{k}^{(3)}$. Plugging all derived results into (40) and applying the property of the generalized Rayleigh quotient, we can obtain the closed-form results in Theorem 4.